# Phase diagram and universal scaling regimes of two-dimensional exciton–polariton Bose–Einstein condensates

### v1

@@ I Introduction
Several universality classes have been identified in non-equilibrium critical phenomena. Among the earliest are the directed percolation universality class in the context of reaction-diffusion processes [1] or the Edwards-Wilkinson (EW) one in the context of stochastic interface growth [2]. The latter leads to non-anomalous exponents and a Gaussian distribution of fluctuations, as in thermal equilibrium, but it plays a key role also in non-equilibrium dynamics such as the Kuramoto-Sivashinsky equation [3, 4] or the deterministic complex Ginzburg-Landau equation [5] when the system has a finite size [6]. The EW universality class was also shown to emerge in exciton-polariton condensates [7, 8, 9]
Another celebrated universality class in non-equilibrium systems is the Kardar–Parisi–Zhang (KPZ) one, originally evidenced in randomly stirred fluids [10] and in the kinetic roughening of growing interfaces [11]. The corresponding dynamics was successively modeled by a stochastic continuous non-linear equation proposed by Kardar, Parisi, and Zhang [11]. KPZ processes display anomalous critical exponents and non-Gaussian distributions of the fluctuations [12, 13]. Far beyond its initial scope of application, the scale-invariant properties defining the KPZ universality class are nowadays observed in a wide variety of physical non-equlibrium phenomena, ranging from biophysics – cancer cells proliferation [14], to quantum physics: integrable spin chains [15, 16], density fluctuations of an Anderson localized wavepacket [17], as well as in phase dynamics of exciton-polariton Bose-Einstein condensates (BEC) [18].
@@
In two-dimensional (2d) systems, even more scenarii for EP universal behavior have been put forward. As specific of the 2d geometry, a phase transition analoguous to the Berezinskii–Kosterlitz–Thouless (BKT) one [27, 28] may arise, between a quasi-ordered phase of bounded vortex-antivortex pairs, and a disordered phase of freely moving vortices. These phases are characterized respectively by algebraic and exponential decays of the condensate spatio-temporal correlation function. This algebraic decay was reported in experiments in quasi-equilibrium EP polaritons [29]. In non-equilibrium conditions, the binding-unbinding transition of vortex-antivortex pairs has been studied theoretically building a non-linear electrodynamics analogy between vortices and charges [30, 31]. This has led to an estimate of the lengthscale $L_{v}$ separating attractive from repulsive interactions between vortex-antivortex pairs, as a function of the microscopic parameters [31]. This behavior was confirmed numerically [32, 33], and microscopic parameters such as the driving intensity are known to play a crucial role in this transition [29, 34, 35]. Numerical studies of EP BECs close to the condensation threshold also report power-law behavior of correlations with an exponent different from the equilibrium one [34].
A KPZ regime may also be found in 2d exciton-polaritons, although the conditions for achieving it are not trivial. In isotropic homogeneous incoherently pumped EP BECs, the condensate is dominated by vortex proliferation which destroys the KPZ phase [36, 37]. Numerical simulations have displayed the emergence of a KPZ regime in idealized conditions, i.e. in the low noise limit and in the absence of polariton-polariton interactions [8], or in the presence of a lattice, which introduces a cutoff length to vortices with smaller cores thereby taming the proliferation of both spatial and space-time vortices [23]. The KPZ regime has also been predicted to occur in anisotropic condensates [36, 38, 39, 40]. Other pumping schemes such as coherent driving [37] or the optical parametric oscillator regime (OPO) [41] also display KPZ features.
@@
We identify as tuning parameter the reservoir-polariton interaction energy, and show that when increasing its value three different regimes are achieved: the Edwards-Wilkinson, Kardar-Parisi-Zhang and vortex phases. We then provide a detailed characterization of these three regimes. In particular, our analysis highlitghts the possibility of experimental observation of the EW regime, characterized by power-law decay of space and time correlations, for which we predict specific ratios of spatial and temporal power-law exponents, which are different from the equilibrium BKT case. The EW regime is the precursor of the KPZ phase, as it emerges at small non-linearities on time and length scales where KPZ fluctuations do not yet build up [44]. Our analysis hence suggests a path towards the experimental observation of KPZ phase. Finally, we comparatively show the properties of the vortex phase highilighting the important differences with respect to the other two phases.
@@ II Model for the condensate dynamics
where $\mathcal{F}^{-1}$ denotes the inverse Fourier transform, with $\epsilon_{\hat{\bm{k}}}$ the dispersion relation on the lower-polariton (LP) branch. Two-dimensional polariton condensates can be realized on honeycomb lattices [42], where the LP branch takes on the characteristic shape of the graphene dispersion relation. Populating the condensate at the $\Gamma$ point of the Brillouin zone, around $\hat{\bm{k}}=\bm{0}$, the condensate dispersion relation is well approximated by the parabola $\epsilon_{\hat{\bm{k}}}=\hbar^{2}\bm{k}^{2}/(2m)$ where $m$ is the (negative) polaritonic effective mass. It can be measured from the full imaging of the LP branch before the condensation threshold. Similarly, the momentum-dependent polariton loss rate $\gamma_{\ell}(\hat{\bm{k}})$ is written within the quadratic approximation $\gamma_{\ell}(\hat{\bm{k}})=\gamma_{0}+\gamma_{2}\bm{k}^{2}$, where the pararmeters $\gamma_{0,2}$ are accessed from the condensate inhomogeneous spectral linewidth [24]. The interaction strength between polaritons is denoted by $g$, while $g_{R}$ represents the interaction strength between the exitonic reservoir and polaritons. The complex noise $\xi$ is Gaussian with zero mean $\langle\xi\rangle=0$ and a covariance $\langle\xi(\bm{r},t)\xi^{*}(\bm{r^{\prime}},t^{\prime})\rangle=2\sigma\delta(\bm{r}-\bm{r^{\prime}})\delta(t-t^{\prime})$. The amplitude $\sigma$ originates from repeated gains from the reservoir and dissipation in the environment and reads $\sigma=\frac{Rn_{R}+\gamma_{0}}{4}$ [19, 46].
@@ III.1 KPZ effective dynamics
The above mapping establishes a connection between the effective dynamics of a phase $\theta(\bm{r},t)$ and that of a classical interface of height $h(\bm{r},t)$ stochastically growing in the normal direction to the two-dimensional plane $\bm{r}$. This interface roughens in space and time and is characterized by a dynamical exponent $z$, a roughness exponent $\alpha$, a growth exponent $\beta=\alpha/z$, related in all dimensions by the exact identity $\alpha+z=2$, stemming from the invariance of the KPZ equation (3) under an infinitesimal tilt of the interface [48]. In 1d, the values of the KPZ critical exponents are exactly known $\alpha=3/2$, $z=1/2$, $\beta=1/3$. In 2d, no exact result is available but they have been determined through extensive numerical simulations of systems belonging to it, giving the estimates $\alpha\approx 0.39$, $z\approx 1.61$, $\beta\approx 0.24$ [49, 50, 51, 52, 53, 54, 55, 56, 57].
@@
Beyond its correlation properties, the universal features of the 2d growing interface can be classified to a greater extent from the distribution of its fluctuations, which indicates the existence of universality subclasses [59, 60, 61, 62]. These subclasses are similar to the ones demonstrated for 1d interfaces in Refs. [63, 13, 12]. They correspond to different global geometries of the growth: flat, curved or stationary. They are associated with different probability distributions although sharing the same set of critical exponents. In contrast with the 1d case, where the form of these distributions is known exactly, being respectively the Tracy-Widom GOE, Tracy-Widom GUE, and Baik-Rains for these three subclasses [63, 13, 12], there is no analytical form for the 2d distributions.
@@ III.2 Weakly nonlinear EW regime
The Edwards-Wilkinson equation corresponds to the linear limit of the KPZ equation (3), i.e. $\lambda=0$ [2], from which the EW universality class is defined. Owing from its linearity, the properties of this equation can easily be worked out, and one can obtain in generic spatial dimension $d$ the critical exponents $\alpha_{\rm EW}=\frac{2-d}{2}$, $z_{\rm EW}=2$, $\beta_{\rm EW}=\frac{2-d}{4}$ as well as its correlation functions [44].
In both one and two dimensions, the EW fixed point of Eq. (3) is IR unstable when the nonlinearity $g_{\rm KPZ}$ is non-zero, and the renormalization group flow eventually reaches the KPZ fixed point. This means that the 2d interface always roughens. As a consequence, when $\lambda\ll 1$, the scaling behavior of $C_{hh}(r,t)$ can be compatible with EW properties at small scales and at the early stages of the growth, but displays KPZ properties beyond some spatiotemporal scale which depends on $1/g_{\rm KPZ}$ [44]. This scenario was confirmed from numerical simulations of lattice models of interface growth, from the gradual change of the exponents with the nonlinearity parameter [55]. Finite-size effects have also been investigated for interfaces within the EW universality class, and for the crossover between the EW and the KPZ class [55, 64, 65].
@@ III.3 Defects-dominated regime
A major difference between the usual KPZ equation [11] and the one obtained in Eq. (3) lies in the compactness of the phase $\theta(\bm{r},t)$, which is defined on the support $\left[-\pi,\pi\right[$, whereas the interface $h(\bm{r},t)$ grows unbounded in $\mathbb{R}$. This specific feature has deep consequences on the growth of the phase interface $\theta(\bm{r},t)$, as it renders possible the formation and proliferation of topological defects. Their nucleation has been studied building a dual electrodynamics theory for the compact version of the KPZ equation [30], and it was shown from a perturbative treatment in the nonlinearity $\lambda$ that vortex-antivortex pairs should undergo a binding-unbinding mechanism, separating an algebraic quasi-ordered phase from a disordered one [31, 38]. This mechanism is characterized by a lengthscale $L_{v}$ representing the typical defect interdistance, that depends on the effective parameters $\nu$, $\lambda$ and $D$ of the KPZ equation. Numerical simulations of the compact KPZ equation indicate that defect nucleation is hindered for systems smaller than $L_{v}$ [39].
@@ IV.1 Numerical simulations
We solve numerically the gGPE (1) with periodic boundary conditions by propagating the initial condition in time with a symmetrized split-step method [77]. We implement the stochastic part of the dynamics by adding to the wave function $\psi(\bm{r},t)$, at each time step and for each spatial point, a complex random number generated by a Euler-Maruyama algorithm, reproducing the statistical properties of the Gaussian noise $\xi$. We solve the reservoir dynamics (2) by using a first order Euler scheme.
The parameters of the simulation are chosen to reproduce typical experimental conditions in GaAs microcavities [24]: $\hbar\gamma_{0}=197.5{\rm\mu eV}$, $\hbar\gamma_{2}=1.77\times 10^{4}{\rm\mu eV.\mu m^{2}}$, $\gamma_{R}=0.45\gamma_{0}$, $R=57.3\times 10^{-3}{\rm\mu m.ps^{-1}}$, $P=2P_{\rm th}$. The polariton-polariton interaction strength $g$ is set to zero, and the interaction between the exciton reservoir and polaritons $g_{R}$ is chosen from the value of the blueshift at threshold $\mu_{\rm th}$ and the steady-state reservoir density, $g_{R}=\frac{1}{2}\mu_{\rm th}/n_{R0}$. The blueshift at threshold $\mu_{\rm th}$ is varied from $50{\rm\mu eV}$ to $700{\rm\mu eV}$ in order to access the different universal regimes presented in Sec. III. The polariton mass is chosen negative $m=-4\times 10^{-5}m_{e}$ so as to counterbalance the sign of the effective interactions ($g_{e}<0$) and preserve the condensate stability [42]. The noise amplitude is approximated in the following by $\sigma\approx Rn_{R}/2$.
@@ IV.2 Identification of the three regimes
In the following, we show that the three regimes presented in Sec. III are accessible by tuning the nonlinear parameter $g_{\rm KPZ}=\frac{\lambda^{2}D}{\nu^{3}}$, whose dependence on the microscopic parameters is known from Eqs. (4)–(6). This effective nonlinearity can be tuned over a broad range of values by changing the blueshift at threshold $\mu_{\rm th}$, which controls the effective polariton interaction strength $g_{e}$. It can be modified through the excitonic fraction in the polaritonic state [43], and it thus provides a natural candidate to explore the universal properties of the 2d condensate.
@@
We solve the condensate dynamics Eqs. (1)–(2) for many values of $\mu_{\rm th}$, and indeed identify three distinct regimes, which correspond to the EW, the KPZ and the vortex regimes of Sec. III. We first provide a qualitative picture, by displaying in Fig. 1 typical density and phase maps, which are snapshots at an arbitrary instant in the steady state, for each of these regimes. The density (left column) appears almost identical in both the EW (upper row) and the KPZ (middle row) regimes. In the vortex phase (bottom row), the condensate density visibly spreads out toward lower values, and interacting vortices are discernible. Regarding the phase (right column), a clear difference is however noticeable between the EW and KPZ regimes from the spread of the phase values. In the vortex regime, the phase explores entirely its compact support and phase defects are visible at spatial points mixing all the colors together, and coinciding with holes in the density map, indicating a strong correlation between the density and phase fields.

### v3

@@ I Introduction
Several universality classes have been identified in non-equilibrium critical phenomena. Among the earliest is the directed percolation universality class in the context of reaction-diffusion processes [1]. Another celebrated universality class in non-equilibrium systems is the Kardar–Parisi–Zhang (KPZ) one, originally evidenced in randomly stirred fluids [10] and in the kinetic roughening of growing interfaces [11]. The corresponding dynamics was successively modeled by a stochastic continuous non-linear equation proposed by Kardar, Parisi, and Zhang [11]. KPZ processes display anomalous critical exponents and non-Gaussian distributions of the fluctuations [12, 13].
In the context of stochastic interface growth, another important universality class is the Edwards-Wilkinson (EW) one [2]. It corresponds to a linear equilibrium dynamics, leading to non-anomalous exponents and a Gaussian distribution of fluctuations. However, it also emerges in non-equilibrium dynamics such as the KPZ equation, the Kuramoto-Sivashinsky equation [3, 4] or the deterministic complex Ginzburg-Landau equation [5] at early times, or when the system has a finite size [6].
Far beyond its initial scope of application, the scale-invariant properties defining the KPZ universality class are nowadays observed in a wide variety of physical non-equlibrium phenomena, ranging from biophysics – cancer cells proliferation [14], to quantum physics: integrable spin chains [15, 16], density fluctuations of an Anderson localized wavepacket [17], as well as in phase dynamics of exciton-polariton Bose-Einstein condensates (BEC) [18]. Note that the EW universality class was also shown to emerge in such condensates [7, 8, 9]
@@
In two-dimensional (2d) systems, even more scenarii for EP universal behavior have been put forward. As specific of the 2d geometry, a phase transition analogous to the Berezinskii–Kosterlitz–Thouless (BKT) one [27, 28] may arise, between a quasi-ordered phase of bounded vortex-antivortex pairs, and a disordered phase of freely moving vortices. These phases are characterized respectively by algebraic and exponential decays of the condensate spatio-temporal correlation function. The algebraic decay was reported in experiments in quasi-equilibrium EP polaritons [29]. In non-equilibrium conditions, the binding-unbinding transition of vortex-antivortex pairs has been studied theoretically building a non-linear electrodynamics analogy between vortices and charges [30, 31]. This has led to an estimate of the lengthscale $L_{v}$ separating attractive from repulsive interactions between vortex-antivortex pairs, as a function of the microscopic parameters [31]. This behavior was confirmed numerically [32, 33], and microscopic parameters such as the driving intensity are known to play a crucial role in this transition [29, 34, 35]. Numerical studies of EP BECs close to the condensation threshold also report power-law behavior of correlations with an exponent different from the equilibrium one [34].
A KPZ regime may also be found in 2d exciton-polaritons, although the conditions for achieving it are not trivial. In isotropic homogeneous incoherently pumped EP BECs, the condensate tends to be dominated by vortex proliferation which destroys the KPZ phase [36, 37]. However, numerical simulations have displayed the emergence of a KPZ regime in idealized conditions, i.e. in the low noise limit and in the absence of polariton-polariton interactions [8], or in the presence of a lattice, which introduces a cutoff length to vortices with smaller cores thereby taming the proliferation of both spatial and space-time vortices [38]. The KPZ regime has also been predicted to occur in anisotropic condensates [36, 39, 40, 41]. Other pumping schemes such as coherent driving [37] or the optical parametric oscillator regime (OPO) [42] also display KPZ features.
@@
We identify as tuning parameter the reservoir-polariton interaction energy, and show that when increasing its value three different regimes are achieved: the Edwards-Wilkinson, Kardar-Parisi-Zhang and vortex phases. We then provide a detailed characterization of these three regimes. In particular, our analysis highlitghts the possibility of experimental observation of the EW regime, characterized by power-law decay of space and time correlations, for which we predict specific ratios of spatial and temporal power-law exponents, which are different from the equilibrium BKT case. The EW regime is the precursor of the KPZ phase, as it emerges at small non-linearities on time and length scales where KPZ fluctuations do not yet build up [45]. Our analysis hence suggests a path towards the experimental observation of KPZ phase. Finally, we comparatively show the properties of the vortex phase highlighting the important differences with respect to the other two phases.
@@ II Model for the condensate dynamics
[…]
@@ III.1 KPZ effective dynamics
The above mapping establishes a connection between the effective dynamics of a phase $\theta(\bm{r},t)$ and that of a classical interface of height $h(\bm{r},t)$ stochastically growing in the normal direction to the two-dimensional plane $\bm{r}$. This interface roughens and becomes scale-invariant, characterized by a dynamical exponent $z$, a roughness exponent $\alpha$, a growth exponent $\beta=\alpha/z$, related in all dimensions by the exact identity $\alpha+z=2$, stemming from the invariance of the KPZ equation (3) under an infinitesimal tilt of the interface [49]. In 1d, the values of the KPZ critical exponents are exactly known $\alpha=3/2$, $z=1/2$, $\beta=1/3$. In 2d, no exact result is available but they have been determined through extensive numerical simulations of systems belonging to the KPZ class, giving the estimates $\alpha\approx 0.39$, $z\approx 1.61$, $\beta\approx 0.24$ [50, 51, 52, 53, 54, 55, 56, 57, 58].
@@
Beyond its correlation properties, the universal features of the 2d growing interface can be classified to a greater extent from the distribution of its fluctuations, which indicates the existence of universality subclasses [60, 61, 62, 63]. These subclasses are similar to the ones demonstrated for 1d interfaces in Refs. [64, 13, 12]. They correspond to different global geometries of the growth: flat, curved or stationary. They are associated with different probability distributions although sharing the same set of critical exponents. In contrast with the 1d case, where the forms of these distributions are known exactly, being respectively the Tracy-Widom GOE, Tracy-Widom GUE, and Baik-Rains for the three subclasses [64, 13, 12], there is no analytical form for the 2d distributions.
@@ III.2 Weakly nonlinear EW regime
The Edwards-Wilkinson equation corresponds to the linear limit of the KPZ equation (3), i.e. setting $\lambda=0$ [2], from which the EW universality class is defined. The EW equation satisfies the classical limit of the fluctuation-dissipation theorem, and it describes thermal-like fluctuations around a stationary interface. Owing from its linearity, the properties of this equation can easily be worked out, and one can obtain in generic spatial dimension $d$ the critical exponents $\alpha_{\rm EW}=\frac{2-d}{2}$, $\beta_{\rm EW}=\frac{2-d}{4}$, $z_{\rm EW}=2$ which corresponds to a diffusive behavior, as well as its correlation functions [45].
In both one and two dimensions, the EW fixed point of Eq. (3) is IR unstable when the nonlinearity $g_{\rm KPZ}$ is non-zero, and the renormalization group flow eventually reaches the KPZ fixed point. This means that the 2d interface always roughens. As a consequence, when $\lambda\ll 1$, and for a smooth initial condition, the scaling behavior of $C_{hh}(r,t)$ can be compatible with EW properties at small scales and at the early stages of the growth, but displays KPZ properties beyond some spatiotemporal scale which depends on $1/g_{\rm KPZ}$ [45]. This scenario was confirmed from numerical simulations of lattice models of interface growth, from the gradual change of the exponents with the nonlinearity parameter [56]. Finite-size effects have also been investigated for interfaces within the EW universality class, and for the crossover between the EW and the KPZ class [56, 65, 66].
@@ III.3 Defects-dominated regime
A major difference between the usual KPZ equation [11] and the one obtained in Eq. (3) lies in the compactness of the phase $\theta(\bm{r},t)$, which is defined on the support $\left[-\pi,\pi\right[$, whereas the interface $h(\bm{r},t)$ grows unbounded in $\mathbb{R}$. This specific feature has deep consequences on the growth of the phase $\theta(\bm{r},t)$, as it renders possible the formation and proliferation of topological defects. Their nucleation has been studied building a dual electrodynamics theory for the compact version of the KPZ equation [30], and it was shown from a perturbative treatment in the nonlinearity $\lambda$ that vortex-antivortex pairs should undergo a binding-unbinding mechanism, separating an algebraic quasi-ordered phase from a disordered one [31, 39]. This mechanism is characterized by a lengthscale $L_{v}$ representing the typical defect interdistance, that depends on the effective parameters $\nu$, $\lambda$ and $D$ of the KPZ equation. Numerical simulations of the compact KPZ equation indicate that defect nucleation is hindered for systems smaller than $L_{v}$ [40].
@@ IV.1 Numerical simulations
We solve numerically the gGPE (1) with periodic boundary conditions by propagating the initial condition in time with a symmetrized split-step method [77]. We implement the stochastic part of the dynamics by adding to the wave function $\psi(\bm{r},t)$, at each time step and for each spatial point, a complex random number generated by a Euler-Maruyama algorithm, reproducing the statistical properties of the Gaussian noise $\xi$. We solve the reservoir dynamics (2) by using a first order Euler scheme. The codes and data produced are available in Ref. [78].
The parameters of the simulation are chosen to reproduce typical experimental conditions in GaAs microcavities [24]: $\hbar\gamma_{0}=197.5{\rm\mu eV}$, $\hbar\gamma_{2}=1.77\times 10^{4}{\rm\mu eV.\mu m^{2}}$, $\gamma_{R}=0.45\gamma_{0}$, $R=57.3\times 10^{-3}{\rm\mu m.ps^{-1}}$, $P=2P_{\rm th}$. The polariton-polariton interaction strength $g$ is set to zero, and the interaction between the exciton reservoir and polaritons $g_{R}$ is chosen from the value of the blueshift at threshold $\mu_{\rm th}$ and the steady-state reservoir density, $g_{R}=\frac{1}{2}\mu_{\rm th}/n_{R0}$. The blueshift at threshold $\mu_{\rm th}$ is varied from $50{\rm\mu eV}$ to $700{\rm\mu eV}$ in order to access the different universal regimes presented in Sec. III.
Note that, in previous works where KPZ universality was evidenced in 2D condensates [8, 38], the condensate dynamics was studied in the adiabatic approximation, neglecting the reservoir dynamics. Moreover, in Ref. [8], the parameter $\gamma_{2}$, which plays a crucial role in the stability of the condensate, was set to zero, and the study was performed in the low noise limit. In Ref. [38], $\gamma_{2}\neq 0$ and the noise was nominal, but the interactions were set to zero. Here, we consider non-zero interactions, and the full non-adiabatic description. Beside being more realistic from an experimental viewpoint, setting a non-zero interaction strength $g_{R}$ dramatically affects the phase diagram of positive mass polariton condensates, creating unstable modes even in the adiabatic approximation of the reservoir. The polariton mass is thus chosen negative $m=-4\times 10^{-5}m_{e}$ so as to counterbalance the sign of the effective interactions ($g_{e}<0$) and stabilize the condensate [43]. Moreover, non-adiabatic effects in the exciton reservoir can give rise to yet other unstable modes or to temporal modulation of the first order correlation function [26]. Including all these elements in our numerics allows us to explore a rich and diverse phase diagram, while addressing the question of the accessibility of KPZ features in the most general setting. The noise amplitude is approximated in the following by $\sigma\approx Rn_{R}/2$.
@@ IV.2 Identification of the three regimes
In the following, we show that the three regimes presented in Sec. III are accessible by tuning the nonlinear parameter $g_{\rm KPZ}=\frac{\lambda^{2}D}{\nu^{3}}$, whose dependence on the microscopic parameters is known from Eqs. (4)–(6). This effective nonlinearity can be tuned over a broad range of values by changing the blueshift at threshold $\mu_{\rm th}$, which controls the effective polariton interaction strength $g_{e}$. Experimentally, the blueshift at threshold can be modified by changing the spectral detuning between the exciton resonance and the cavity mode, thus tuning the excitonic fraction of the polariton state [44, 79]. It thus provides a natural candidate to explore the universal properties of the 2d condensate. Note that similar results could be obtained varying the pumping rate $p$ instead.
@@
We solve the condensate dynamics Eqs. (1)–(2) for many values of $\mu_{\rm th}$, and indeed identify three distinct regimes, which correspond to the EW, the KPZ and the vortex regimes of Sec. III. We first provide a qualitative picture, by displaying in Fig. 1 typical density and phase maps, which are snapshots at an arbitrary instant in the steady state, for each of these regimes. The density (left column) appears almost identical in both the EW (upper row) and the KPZ (middle row) regimes. In the vortex phase (bottom row), the condensate density visibly spreads out toward lower values, and interacting vortices are discernible. Regarding the phase (right column), a clear difference is however noticeable between the EW and KPZ regimes from the spread of the phase values. In the vortex regime, the phase explores entirely its compact support and phase defects are visible at spatial points mixing all the colors together, and coinciding with holes in the density map, indicating a strong correlation between the density and phase fields. We have checked that the phase diagram remains similar decreasing the pumping rate down to $p\simeq 1.5$, below which the proliferation of space-time defects also affects the condensate dynamics [25, 26].